A beaker with 110 mL of an acetic acid buffer with a pH of 5.00 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 M. A student adds 8.80 mL of a 0.400 MM HCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.760.
Answer #1

Answer

∆pH = -0.58

Explanation

Henderson - Hasselbalch equation is

pH = pKa + log([A-]/[HA])

5.00 = 4.760 + ([CH3COO-]/[CH3COOH])

log[CH3COO-] /[CH3COOH] = 0.24

[CH3COO-] / [CH3COOH]= 1.738

[CH3COO-] = 1.738[CH3COOH]

[CH3COO-] + [CH3COOH] = 0.100M

(1.738 × [CH3COOH]) + [CH3COOH] = 0.100M

2.738 × [CH3COOH] = 0.100M

[CH3COOH] = 0.03652M

[CH3COO-] = 0.100M - 0.03652M = 0.06348M

moles of CH3COOH = ( 0.03652mol/1000ml) × 110ml = 0.0040172mol

moles of CH3COO- = ( 0.06348mol/1000ml) × 1!0ml = 0.0069828mol

HCl react with conjucate base CH3COO-

CH3COO- + H+ --------> CH3COOH

moles of HCl added = ( 0.400mol/1000ml) × 8.80mln= 0.00352 mol

after addition of HCl

moles of CH3COOH = 0.0040172mol + 0.00352mol = 0.0075372mol

moles of CH3COO- = 0.0069828mol - 0.00352mol = 0.0034627mol

Total volume = 110ml + 8.80ml = 118.80ml

[CH3COOH] = (0.0075372mol/118.80ml) ×1000ml = 0.06344M

[CH3COO-] = (0.0034627mol/118.80ml) × 1000ml = 0.02915M

Applying Henderson - Hasselbalch equation

pH = 4.76 + log( 0.02915M/0.06344M)

pH = 4.76 - 0.34

pH = 4.42

∆pH = 4.42 - 5.00

∆pH = - 0.58